OpenStudy (anonymous):

GIVING MEDAL TO BEST ANSWER! I'D APPRECIATE HELP! A tank holds 400 gallons of water. A crack forms and 20 gallons of water leak out every hour. Write and simplify an expression to find how much water is in the tank after 12 hours.

OpenStudy (anonymous):

The volume of water in the tank decreases by 20 gallons per minute, so you can set up a pretty simple differential equation relating the change in volume \(V\): \[\frac{dV}{dt}=-20\] Solving yields \(V=-20t+C\). You know that at time \(t=0\), you have a volume of 400 gallons (assuming the tank to be full initially). This means you have \[400=-20(0)+C~~\implies~~C=400\] and so the volume at any time \(t\ge0\) is \[V=-20t+400\]